OpenStudy (anonymous):

how many moles of methane are produced when 48.1 moles of carbon dioxide react with excess hydrogen gas?

OpenStudy (anonymous):

(*48.1 grams) and also given the equation \[C0_{2}+4H _{2}rightarrowCH4+2H _{2}0\]

OpenStudy (abmon98):

Use Equation Number of Moles=Mass(g)/Molar Mass(g/mol) Use each element in CO2 atomic weight C:12 g O:16 g Molar Mass=(12*1)+(16*2)=44 g

OpenStudy (anonymous):

thank you!!

OpenStudy (abmon98):

you got it?, your most welcome :)